Help with my homework question please.

11. Calculate the surface integral, (16** +°)e=ds, where S is the cylinder x² + y2 = 9 for 0 <<1.
Answer #1

+ т (1,1) = 3сол) 7+ 3 sin 9г 17 r 1- 3Co49, 9- ? { in 9, 37,7)= z 8 = 4 6 * с = 0-5 -1 +0°0,4 6 =ght at z= (х 16 Guose = h G0 49 1 2) 스Z스 단 tt IXR 깁크 C Z. do - 3sing T +309 0 안시 55 ㅋ = 6 이 -3sing 3Cos9. - Fro - 2010) - Cottoo) + R (0)daar =-3coso 7 -3 sino Х 29 m - 19 = 3 S 6x² + y²) e?. cos? + r sin?) e x3. JA isy??3.dz do = o con 2n e 02711 dad 0124.7.786 LATE 27 -2 27 ( e 2) - no do 201 27 (e!_c 2 . ( o -) = -24 - 271 TIEZ 597 (1) 34:137 I 107.23 Ang